OpenStudy (anonymous):

If the sum of three consecutive integers is 57, what are the integers?

OpenStudy (amistre64):

yay!! one I can solve :)

OpenStudy (amistre64):

lets start by saying the the first number in our little play is an "x" and we want to find out what its value is... the next 2 numbers from x are: x+1 and x+2 we know that (x) + (x+1) + (x+2) = 57 so lets solve it ;)

OpenStudy (amistre64):

3x+3 = 57 3x = 54 x = 54/3 x = 18 the sequence is: 18,19,20

OpenStudy (anonymous):

thanks!